sapana is six years younger than kristen. six years ago, kristen was three times as old as sapana was than. how old was sapana six years ago?
Answer by stanbon(75887)

sapana is six years younger than kristen. six years ago, kristen was three times as old as sapana was than. how old was sapana six years ago?
----
Equations:
s = k-6
k-6 = 3(s-6)
----
Substitute for "s" and solve for k:
k-6 = 3(k-6-6)
k-6 = 3k - 36
2k = 30
k = 15 (kristen's age now)
---
Solve for "s":
s = k-6
s = 15-6 = 9 (sapana's age now)
s-6 = 9-6 = 3 (sapana's age 6 years ago)
=============================================
Cheers,
stan H.
----